# SGDFuse: SAM-Guided Diffusion for High-Fidelity Infrared and Visible Image Fusion

### v1

@@ Abstract
Index Terms:
Image Fusion, Diffusion Model, Segment Anything Model (SAM), Multimodal Information
I Introduction
Infrared and visible image fusion is an important research direction in the field of computer vision [1], [2]. This technology aims to efficiently combine the thermal radiation information of infrared images with the rich texture details of visible light images, to generate an image that is more comprehensive in information and higher in perceptual clarity [3], [4]. Because infrared images can effectively overcome adverse conditions such as smoke and low
@@ I Introduction
light, while visible light images possess the advantages of high resolution and high contrast, the features of the two are highly complementary. This fusion not only significantly improves the environmental perception and target recognition ability in key scenarios such as autonomous driving [5], military reconnaissance, and medical image analysis [6], but also effectively enhances the performance of downstream vision tasks [7]. Therefore, how to effectively achieve fusion to balance the final visual quality and task requirements, is an urgent challenge currently being faced.
In pursuit of this objective, a plethora of techniques have been proposed over the past few decades [8]. Conventional fusion methods are largely predicated on handcrafted feature extraction and fusion rules—such as sparse representation [9], multi-scale transforms [10], and subspace reconstruction [11]—which offer a degree of interpretability and robustness. However, these approaches face inherent limitations in modeling the high-order semantic relationships and non-linear complementarities between modalities, thereby restricting their efficacy in complex scenes [12]. With the advent and rapid proliferation of Deep Learning (DL), the research landscape for IVIF has undergone a paradigm shift. This has yielded substantial advancements in feature extraction, fusion, and reconstruction, rapidly establishing the field as a prominent area of research. For instance, Li et al. introduced DenseFuse [13], which leverages an encoder-decoder architecture to extract multi-level features; PIAFusion [14], proposed by Tang et al., employs a progressive fusion strategy to handle extreme variations in illumination; DATFuse [15] further incorporates a dual-attention Transformer to model long-range contextual information; and more recently, Text-DiFuse [16] achieves controllable fusion and enhanced robustness against degradation by integrating text modulation with diffusion-based generation.
Despite the advancements in feature extraction and fusion strategies for infrared and visible images, existing methods universally overlook a central issue: the lack of deep semantic understanding. Ideal fusion should transcend the mere superposition of pixel information, evolving into a process of selective enhancement guided by semantic importance—meaning the model must discern the target from the background. However, current mainstream approaches remain confined to low-level visual cues like pixel intensity and gradients, rendering them incapable of explicitly modeling high-level semantic regions such as targets and their boundaries. This ”can’t see the forest for the trees” paradigm directly leads to a cascade of problems, including blurred target boundaries, the loss of critical structures, and the erroneous suppression of thermal signatures. Ultimately, this semantic deficiency is the fundamental bottleneck constraining the efficacy of fused images in downstream tasks like object detection and semantic segmentation. As illustrated in Fig. 1, the model’s failure to preserve task-critical target regions severely undermines the practical utility of fusion technology.
@@
Recent advancements in large-scale vision models offer new promise [17]. The Segment Anything Model (SAM) [18] represents a key breakthrough owing to its outstanding semantic segmentation capabilities. The high-quality semantic masks it generates can provide multi-modal images with rich semantic priors (as shown in Fig. 2). This wealth of semantic information can significantly alleviate the ”semantic blindness” inherent in existing methods. It guides the model to more effectively preserve the thermal radiation features and visible details of critical targets, thereby achieving a more semantically consistent fused representation. On the other hand, image fusion, as an image generation task, is typically accompanied by processes of information compression and modal reconstruction [19], [20]. Although this process can enhance visual consistency, it often leads to a degradation in detail quality, manifesting as blurred textures, indistinct boundaries, and visual artifacts. Such degradation not only diminishes the perceptual quality of the fused image but also significantly impairs the performance of downstream tasks that rely on structural information, such as object detection and semantic segmentation. Diffusion models, by virtue of their iterative denoising generation mechanism, can stably produce high-fidelity, artifact-free images, making them
@@
exceptionally well-suited for the demands of IVIF, namely detail restoration and modal balance. More importantly, when integrated with high-quality semantic priors, diffusion models can perform semantically-aware conditional generation. This enables the fusion process to be more targeted in its representation of structures, textures, and target regions, significantly enhancing its adaptability for downstream tasks.
@@
II Related work
II-A Deep Multimodal Image Fusion
@@
Autoencoder (AE)-based methods, such as DenseFuse [13] and NestFuse [22], utilize an encoder-decoder architecture for multi-scale feature extraction and fusion. While advanced variants like SEDRFuse [23] introduce refinements using residual and attention mechanisms, their fundamental reliance on pixel-level loss restricts the capture of deep semantics, often leading to a loss of critical contextual information.
@@
Transformer-based methods provide a new fusion paradigm, leveraging long-range dependency modeling for superior cross-domain integration. Various works have employed Transformers to enhance inter-modal dependencies, suppress redundancy, and integrate complementary features [26], [27]. Building on this foundation, research has extended into more specialized and cutting-edge applications. Li et al. [28] proposed a U-Net-based framework, pioneering the exploration of adversarial robustness in fusion tasks. To achieve precise semantic guidance, Liu et al. [29] pioneered the application of prompt learning to infrared and visible image fusion, with achieving sota performance.
@@
II-B SAM and Its Applications
The Segment Anything Model (SAM) [18], as a powerful pretrained model, has become a current research hotspot in the field of image processing due to its zero-shot generalization capability in image segmentation and object detection tasks [32]. SAM can provide high-precision semantic masks across various vision tasks and has been widely applied in low-level vision tasks such as deblurring [33], dehazing [34], and low-light image enhancement [35], as well as high-level tasks like object detection [36], demonstrating outstanding performance. By deeply understanding image content, SAM effectively captures fine details and accurately delineates target regions. Therefore, SAM is particularly suitable for infrared and visible image fusion tasks, we leverage SAM-generated segmentation masks as explicit, high-level semantic priors to guide the fusion process.
II-C Diffusion Models
Diffusion models have emerged as a powerful class of deep generative models, achieving state-of-the-art results in image generation, restoration, super-resolution, and image-to-image translation tasks [7], [37], [38]. Beyond generative applications, their learned representations have also shown potential in discriminative tasks such as classification, segmentation, and object detection. A diffusion model consists of two stages: a forward process that gradually adds Gaussian noise to the input data, and a reverse process that iteratively denoises the corrupted data by predicting and removing noise, ultimately recovering high-quality samples [38], [39].
The advantages of diffusion models lie in their ability to generate high-quality and diverse samples while effectively capturing deep underlying features of the data, which leads to outstanding performance in image generation tasks [40]. With the continuous advancement of diffusion models across various fields, particularly in image generation, they have begun to surpass the long-standing dominance of generative adversarial networks (GANs). In the context of infrared and visible image fusion, diffusion models provide powerful support for generating high-quality fused images.
@@ II-C Diffusion Models
III Methodology
III-A Motivation
Existing IVIF methods face a core bottleneck in semantic understanding, as they treat fusion as a mathematical reorganization of low-level visual features (e.g., pixel intensity, gradients).
@@ III-A Motivation
This failure to semantically distinguish between critical targets and background textures acts as a fundamental obstacle, severely limiting the potential of the fused images in downstream tasks like object detection and semantic segmentation.
Our primary and core motivation is to inject explicit semantic understanding capabilities into the fusion process. We are no longer satisfied with the indiscriminate processing of low-level features by existing methods, but rather seek a novel mechanism that enables the model to truly ”understand” the image and distinguish ”target” from ”background”. For this, we introduce SAM. The high-quality semantic masks it generates can provide the fusion process with unprecedented, strong semantic priors concerning scene structure and object boundaries, which becomes our ideal tool for solving the core problem of ”semantic blindness”. However, high-level semantic guidance alone is insufficient. To accurately translate this precise semantic intent—for instance, ’to preserve the thermal features of a target while enhancing the texture of the background’—into a high-quality, pixel-level output, one must rely on an equally powerful high-fidelity generative framework. Traditional reconstruction methods are prone to generating artifacts and detail loss when handling such complex conditions, which severely undermines the efficacy of semantic guidance. Therefore, our second motivation is to employ a Diffusion Model, renowned for its superior generative capabilities, as the backbone for image reconstruction. This ensures that the high-level semantic commands can be rendered losslessly and with maximum fidelity.
@@ III-B Network Architecture
III-C Multi-Scale Feature Enhancement Module (MSFEM)
@@
III-D Semantic-Aware Fusion Optimization Driven by Conditional Diffusion
To enhance structural refinement and semantic consistency of the fused image, this work introduces an optimization mechanism based on a conditional diffusion model in the second stage. Unlike existing methods that directly concatenate the original infrared and visible images, our approach concatenates the initial fused image $F_{1}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 3}$ generated in the first stage with two semantic masks $M_{ir}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 1}$ and $M_{vis}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 1}$ extracted from the infrared and visible images using the Segment Anything Model (SAM), forming a five-channel task-aware input. We employ the Denoising Diffusion Probabilistic Model (DDPM) [41] to model the diffusion generation process of this five-channel input in the latent space. The forward process gradually adds Gaussian noise over T time steps to produce a perturbed data distribution, while the reverse process progressively denoises the input under the guidance of semantic masks to reconstruct a fused image with clear structure and semantic consistency. By modeling the joint distribution of the fused image and semantic priors, the diffusion model learns a semantic-aware generative pathway, thereby further improving fusion quality and adaptability to downstream tasks [42].
III-D1 Forward Diffusion Process
@@
III-D2 Reverse Diffusion Process
@@
III-D3 Loss Function of Diffusion Process
@@
III-D4 Structure of the Denoising Network
@@
III-E Hierarchical Feature Aggregation Head (HFAH)
@@
III-F Loss Function of Training Process
@@
TABLE I: Quantitative comparison on MSRS and M3FD datasets, with red representing the best results, blue representing the second best results.
@@
TABLE II: Quantitative comparison on LLVIP and RoadScene datasets, with red representing the best results, blue representing the second best results.
@@
III-F1 Stage I: Supervision of Preliminary Fusion
$L_{int}^{1}=\frac{1}{HW}\parallel F_{1}-I_{ir}\parallel,$ (14)
The overall loss for the first stage is defined as:
$L_{stage1}=L_{int}^{1}+L_{grad}^{2},$ (15)
III-F2 Stage II: Diffusion Optimization Training
$\mathrm{M=max(M_{ir},M_{vis})},$ (16)
@@ III-F2 Stage II: Diffusion Optimization Training
Finally, the total loss for the second stage is defined as:
IV EXPERIMENTS
IV-A Experimental Settings
IV-A1 Datasets
@@ IV-A3 Training Details
IV-B Generalization Comparison Experiments
@@
IV-B1 Results and Analysis on the MSRS Dataset
As shown in Table I, our method achieves the best performance across five key metrics: EN, SD, SF, VIF, and $Q^{abf}$. The higher EN and MI values indicate better information preservation from both modalities,
@@ IV-B1 Results and Analysis on the MSRS Dataset
while superior SD and SF reflect enhanced contrast and edge sharpness. Notably, the highest $Q^{abf}$ further confirms the effectiveness of our method in maintaining structural consistency and perceptual quality.
IV-B2 Results and Analysis on the M3FD Dataset
IV-B3 Results and Analysis on the LLVIP Dataset
As shown in Table II, our proposed method attains the highest scores across all evaluation metrics except MI, where it ranks second, demonstrating strong competitiveness. These results highlight SGDFuse’s clear superiority in preserving information content, texture details, edge sharpness, and perceptual quality. Notably, leading performance in EN, SD, and SF indicates richer information, enhanced contrast, and improved clarity in the fused images. Furthermore, the superior $Q^{abf}$ and VIF scores confirm that SGDFuse achieves better perceptual consistency and edge alignment, closely aligning with human visual preferences.
@@ IV-B3 Results and Analysis on the LLVIP Dataset
IV-B4 Results and Analysis on the RoadScene Dataset
As shown in Table II, our method attains the best performance on four key metrics: MI, SCD, VIF, and $Q^{abf}$.
The leading MI score evidences effective preservation of complementary information from infrared and visible inputs. Superior SCD highlights strong structural and contour fidelity. Additionally, top VIF and $Q^{abf}$ scores confirm enhanced perceptual quality and edge alignment, demonstrating the method’s robust capability to produce semantically consistent and visually faithful fusion results.

### v2

@@ Abstract
keywords:
Image Fusion, Diffusion Model, Segment Anything Model (SAM), Multimodal Information, Image Enhancement
journal: Nuclear Physics B
1 Introduction
Infrared and visible image fusion is an important research direction in the field of computer vision [1]. This technology aims to efficiently combine the thermal radiation information of infrared images with the rich texture details of visible light images, to generate an image that is more comprehensive in information and higher in perceptual clarity [2]. Because infrared images can effectively overcome adverse conditions such as smoke and low light, while visible light images possess the advantages of high resolution and high contrast, the features of the two are highly complementary. This fusion not only significantly improves the environmental perception and target recognition ability in key scenarios such as autonomous driving [3], military reconnaissance, and medical image analysis [4], but also effectively enhances the performance of downstream vision tasks [5]. Therefore, how to effectively achieve fusion to balance the final visual quality and task requirements, is an urgent challenge currently being faced.
In pursuit of this objective, a plethora of techniques have been proposed over the past few decades [6]. Conventional fusion methods are largely predicated on handcrafted feature extraction and fusion rules—such as multi-scale transforms [7], and subspace reconstruction—which offer a degree of interpretability and robustness. However, these approaches face inherent limitations in modeling the high-order semantic relationships and non-linear complementarities between modalities, thereby restricting their efficacy in complex scenes [8]. With the advent and rapid proliferation of Deep Learning (DL), the research landscape for IVIF has undergone a paradigm shift. This has yielded substantial advancements in feature extraction, fusion, and reconstruction, rapidly establishing the field as a prominent area of research. For instance, Li et al. introduced DenseFuse [9], which leverages an encoder-decoder architecture to extract multi-level features; PIAFusion [10], proposed by Tang et al., employs a progressive fusion strategy to handle extreme variations in illumination; Text-DiFuse [11] achieves controllable fusion and enhanced robustness against degradation by integrating text modulation with diffusion-based generation.
@@ 1 Introduction
Despite advancements in feature extraction and fusion strategies, inherent design limitations in existing methods lead to a core problem: a ’semantic blindness’. For instance, methods based on CNNs are constrained by their local receptive fields, an architectural limitation that hinders their ability to comprehend global scene structures, often resulting in blurred boundaries for large-scale targets or inconsistent background luminance. On the other hand, while GAN-based approaches can produce visually sharp images, their training relies on pixel-level fidelity constraints (e.g., intensity and gradient losses), which struggle to capture complex cross-modal semantic relationships. This not only causes training instability but also frequently leads to the erroneous suppression of salient thermal targets in infrared images. These specific design flaws point to a fundamental issue: the models are unable to semantically differentiate between foreground targets and background textures, thus failing to perform selective information enhancement. This ”can’t see the forest for the trees” paradigm directly causes a cascade of problems, including blurred target boundaries, loss of critical structures, and erroneous suppression of thermal features. Ultimately, this semantic deficiency is the fundamental bottleneck constraining the efficacy of fused images in downstream tasks like object detection and semantic segmentation. As illustrated in Fig. 1, the failure of existing models to preserve task-critical target regions severely undermines the practical utility of fusion technology.
@@
Recent advancements in large-scale vision models offer new promise [12]. The Segment Anything Model (SAM) [13] represents a key breakthrough owing to its outstanding semantic segmentation capabilities. The high-quality semantic masks it generates can provide multi-modal images with rich semantic priors (as shown in Fig. 2). This wealth of semantic information can significantly alleviate the ”semantic blindness” inherent in existing methods. It guides the model to more effectively preserve the thermal radiation features and visible details of critical targets, thereby achieving a more semantically consistent fused representation. On the other hand, image fusion, as an image generation task, is typically accompanied by processes of information compression and modal reconstruction [14]. Although this process can enhance visual consistency, it often leads to a degradation in detail quality, manifesting as blurred textures, indistinct boundaries, and visual artifacts. Such degradation not only diminishes the perceptual quality of the fused image but also significantly impairs the performance of downstream tasks that rely on structural information, such as object detection and semantic segmentation. Diffusion models, by virtue of their
@@
iterative denoising generation mechanism, can stably produce high-fidelity, artifact-free images, making them exceptionally well-suited for the demands of IVIF, namely detail restoration and modal balance. More importantly, when integrated with high-quality semantic priors, diffusion models can perform semantically-aware conditional generation. This enables the fusion process to be more targeted in its representation of structures, textures, and target regions, significantly enhancing its adaptability for downstream tasks.
@@
2 Related work
2.1 Deep Multimodal Image Fusion
@@
Autoencoder (AE)-based methods, such as DenseFuse [9], utilize an encoder-decoder architecture for multi-scale feature extraction and fusion. While advanced variants like SEDRFuse [16] introduce refinements using residual and attention mechanisms, their fundamental reliance on pixel-level loss restricts the capture of deep semantics, often leading to a loss of critical contextual information.
@@
Transformer-based methods provide a new fusion paradigm, leveraging long-range dependency modeling for superior cross-domain integration. Various works have employed Transformers to enhance inter-modal dependencies, suppress redundancy, and integrate complementary features [19]. Building on this foundation, research has extended into more specialized and cutting-edge applications. Li et al. [20] proposed a U-Net-based framework, pioneering the exploration of adversarial robustness in fusion tasks. To achieve precise semantic guidance, Liu et al. [21] pioneered the application of prompt learning to infrared and visible image fusion, with achieving sota performance.
@@
2.2 SAM and Its Applications
The Segment Anything Model (SAM) [13], as a powerful pretrained model, has become a current research hotspot in the field of image processing due to its zero-shot generalization capability in image segmentation and object detection tasks [24]. SAM can provide high-precision semantic masks across various vision tasks and has been widely applied in low-level vision tasks such as dehazing [25], and low-light image enhancement [26], as well as high-level tasks like object detection [27], demonstrating outstanding performance. By deeply understanding image content, SAM effectively captures fine details and accurately delineates target regions. Therefore, SAM is particularly suitable for infrared and visible image fusion tasks, we leverage SAM-generated segmentation masks as explicit, high-level semantic priors to guide the fusion process.
2.3 Diffusion Models
Diffusion models have emerged as a powerful class of deep generative models, achieving state-of-the-art results in image generation, restoration, super-resolution, and image-to-image translation tasks [5], [28]. Beyond generative applications, their learned representations have also shown potential in discriminative tasks such as classification, segmentation, and object detection. A diffusion model consists of two stages: a forward process that gradually adds Gaussian noise to the input data, and a reverse process that iteratively denoises the corrupted data by predicting and removing noise, ultimately recovering high-quality samples [29].
The advantages of diffusion models lie in their ability to generate high-quality and diverse samples while effectively capturing deep underlying features of the data, which leads to outstanding performance in image generation tasks. With the continuous advancement of diffusion models across various fields, particularly in image generation, they have begun to surpass the long-standing dominance of generative adversarial networks (GANs). In the context of infrared and visible image fusion, diffusion models provide powerful support for generating high-quality fused images.
@@ 2.3 Diffusion Models
3 Methodology
3.1 Motivation
Existing IVIF methods face a core bottleneck in semantic understanding, as they treat fusion as a mathematical reorganization of low-level visual features (e.g., pixel intensity, gradients). This failure to semantically distinguish between critical targets and background textures acts as a fundamental obstacle, severely limiting the potential of the fused images in downstream tasks like object detection and semantic segmentation.
Our primary and core motivation is to inject explicit semantic understanding capabilities into the fusion process. We are no longer satisfied with the indiscriminate processing of low-level features by existing methods, but rather seek a novel mechanism that enables the model to truly ”understand” the image and distinguish ”target” from ”background”. For this, we introduce SAM. The high-quality semantic masks it generates can provide the fusion process with unprecedented, strong semantic priors concerning scene structure and object boundaries, which becomes our ideal tool for solving the core problem of ”semantic blindness”. However, high-level semantic guidance alone is insufficient. To accurately translate this precise semantic intent—for instance, ’to preserve the thermal features of a target while enhancing the texture of the background’—into a high-quality, pixel-level output, one must rely on an equally powerful high-fidelity generative framework. Traditional reconstruction methods are prone to generating artifacts and detail loss when handling such complex conditions, which severely undermines the efficacy of semantic guidance. Therefore, our second motivation is to employ a Diffusion Model, renowned for its superior generative capabilities, as the backbone for image reconstruction. This ensures that the high-level semantic commands
@@ 3.1 Motivation
can be rendered losslessly and with maximum fidelity.
@@ 3.2 Network Architecture
3.3 Multi-Scale Feature Enhancement Module (MSFEM)
@@
3.4 Semantic-Aware Fusion Optimization Driven by Conditional Diffusion
To enhance structural refinement and semantic consistency of the fused image, this work introduces an optimization mechanism based on a conditional diffusion model in the second stage. Unlike existing methods that directly concatenate the original infrared and visible images, our approach concatenates the initial fused image $F_{1}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 3}$ generated in the first stage with two semantic masks $M_{ir}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 1}$ and $M_{vis}\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 1}$ extracted from the infrared and visible images using the Segment Anything Model (SAM), forming a five-channel task-aware input. We employ the Denoising Diffusion Probabilistic Model (DDPM) [30] to model the diffusion generation process of this five-channel input in the latent space. The forward process gradually adds Gaussian noise over T time steps to produce a perturbed data distribution, while the reverse process progressively denoises the input under the guidance of semantic masks to reconstruct a fused image with clear structure and semantic consistency. By modeling the joint distribution of the fused image and semantic priors, the diffusion model learns a semantic-aware generative pathway, thereby further improving fusion quality and adaptability to downstream tasks.
@@ 3.4 Semantic-Aware Fusion Optimization Driven by Conditional Diffusion
3.4.1 Forward Diffusion Process
@@
3.4.2 Reverse Diffusion Process
@@
3.4.3 Loss Function of Diffusion Process
@@
3.4.4 Structure of the Denoising Network
@@
3.5 Hierarchical Feature Aggregation Head (HFAH)
@@
3.6 Loss Function of Training Process
@@
3.6.1 Stage I: Supervision of Preliminary Fusion
$L_{int}^{1}=\frac{1}{HW}\parallel F_{1}-I_{ir}\parallel,$ (14)
The overall loss for the first stage is defined as:
$L_{stage1}=L_{int}^{1}+L_{grad}^{2},$ (15)
3.6.2 Stage II: Diffusion Optimization Training
$\mathrm{M=max(M_{ir},M_{vis})},$ (16)
Finally, the total loss for the second stage is defined as:
Table 1: Quantitative comparison on MSRS and M3FD datasets, with red representing the best results, blue representing the second best results.
@@ 3.6.2 Stage II: Diffusion Optimization Training
Table 2: Quantitative comparison on LLVIP and RoadScene datasets, with red representing the best results, blue representing the second best results.
@@
4 EXPERIMENTS
4.1 Experimental Settings
4.1.1 Datasets
@@ 4.1.3 Training Details
4.2 Generalization Comparison Experiments
@@
4.2.1 Results and Analysis on the MSRS Dataset
As shown in Table I, our method achieves the best performance across five key metrics: EN, SD, SF, VIF, and $Q^{abf}$. The higher EN and MI values indicate better information preservation from both modalities, while superior SD and SF reflect enhanced contrast and edge sharpness. Notably, the highest $Q^{abf}$ further confirms the effectiveness of our method in maintaining structural consistency and perceptual quality.
4.2.2 Results and Analysis on the M3FD Dataset
4.2.3 Results and Analysis on the LLVIP Dataset
As shown in Table II, our proposed method attains the highest scores across all evaluation metrics
@@ 4.2.3 Results and Analysis on the LLVIP Dataset
except MI, where it ranks second, demonstrating strong competitiveness. These results highlight SGDFuse’s clear superiority in preserving information content, texture details, edge sharpness, and perceptual quality. Notably, leading performance in EN, SD, and SF indicates richer information, enhanced contrast, and improved clarity in the fused images. Furthermore, the superior $Q^{abf}$ and VIF scores confirm that SGDFuse achieves better perceptual consistency and edge alignment, closely aligning with human visual preferences.
@@
Figure 9: Comparison of different fusion methods in terms of parameter count and computational complexity.
4.2.4 Results and Analysis on the RoadScene Dataset
As shown in Table II, our method attains the best performance on four key metrics: MI, SCD, VIF, and $Q^{abf}$. The leading MI score evidences effective preservation of complementary information from infrared and visible inputs. Superior SCD highlights strong structural and contour fidelity. Additionally, top VIF and $Q^{abf}$ scores confirm enhanced perceptual quality and edge alignment, demonstrating the method’s robust capability to produce semantically consistent and visually faithful fusion results.